OpenStudy (anonymous):

Whats the sum of Radical 18 and Radical 72?

OpenStudy (anonymous):

break the radicals down into prime numbers for instance radical 63 is \[3\sqrt{7}\]

OpenStudy (anonymous):

\[3\sqrt{2} + 6\sqrt{2} = 9\sqrt{2}\]

OpenStudy (anonymous):